There are thirty-three students in the Chess Club. There are five more boys than girls in the club. Write and solve a system of equations to find the number of boys and girls in the Chess Club.
Answer by mananth(12270)

x number of boys
y number of girls
Total attended
x+y = 33 .............1

x-y=5 .............2
Eliminate y

Add the two equations
2 x = 38
/ 2
x = 19
plug value of x in (1)
1 x + 1 y = 33
19 + 1 y = 33
1 y = 33-19
1 y = 14
y = 14
19 number of boys
14 number of girls
m.ananth@hotmail.ca